Opened 9 years ago

Closed 3 years ago

#339 closed flight processing (fixed)

GB09/00, flight day 174/2010, Oxford Parks

Reported by: jb362 Owned by: knpa
Priority: alpha 4 low Milestone: 2010 data processing completion
Component: Archiving Keywords:
Cc: Other processors: emca

Description (last modified by anch)

Data location: ~arsf/workspace/GB08_05-2010_174_Oxford_parks

Data arrived from ARSF via network transfer on 28th June 2010

Scientific objective: Calibration

Priority: alpha 4 low

PI: M Pollard

  • Eagle (20/12/2010)
  • Hawk (20/12/2010)
  • Leica LIDAR (26/08/2010)
  • Photography (26/08/2010)

Change History (33)

comment:1 Changed 9 years ago by jb362

DEM script could not be run on this flight during unpacking. Possibly because computer has old version of GRASS...

comment:2 Changed 9 years ago by benj

  • Description modified (diff)

comment:3 Changed 9 years ago by benj

  • Description modified (diff)

comment:4 Changed 9 years ago by jb362

This directory needs to be backed up in the repository

comment:5 Changed 9 years ago by ella

IPAS navigation processing done.

comment:6 Changed 9 years ago by ella

Applanix navigation processing could not be completed. During the last step in the procedure no sbet file was produced. There is a gap in the flight and we have tried to split the flight but this made no difference.
Files have been copied back to the arsf workspace on samba so if anyone comes to process this project they can pick up where we left off in PosPac.

comment:7 Changed 9 years ago by jb362

Applanix processing completed and copied to workspace

comment:8 Changed 9 years ago by jb362

IPAS processing completed and copied to workspace

comment:9 Changed 9 years ago by anch

  • Owner set to anch
  • Status changed from new to accepted

I have completed the processing of the hyperspectral data, and the project has been submitted for checking.

comment:10 Changed 9 years ago by mark1

OS vectors have arrived and have been unpacked to ~arsf/vectors/GB08_05

comment:11 Changed 9 years ago by anch

SCT Offsets for this project:

EAGLE:
-2: 0.01
-3: 0.00

HAWK:
-2: 0.01
-3: 0.00

comment:12 Changed 9 years ago by emca

Have delivery check the hyperspectral delivery at
~airborne/workspace/GB08_05-2010_174_Oxford_parks/delivery/20100713

Project code is incorrect for this project (this needs to be changed in the directory structure, delivery, logsheet & ticket)
The PI name is missing in the logsheet and the read_me
Need to rename logsheet to correct format
Sample command - need to change the dem name
Lev1 file names should start at 01 i.e. e174011b*
Permissions on lev1 *.bil files needs to be changed to allow readable to all (to avoid problems when copying the data to disk)

comment:13 Changed 9 years ago by anch

  • Summary changed from GB08/05, flight day 174/2010, Oxford Parks to GB09/00, flight day 174/2010, Oxford Parks

comment:14 Changed 9 years ago by anch

I think I have fixed all the above problems including renaming the directory structure.

This probably needs checking, as I may have missed a directory somewhere and therefore not renamed it to GB09_00.

comment:15 Changed 9 years ago by emca

  • Other processors set to emca

Processing LIDAR and photography data

comment:16 Changed 9 years ago by emca

Camera and LiDAR ready for delivery check:

C: ~airborne/workspace/GB09_00-2010_174_Oxford_parks/delivery/20100824
L: ~airborne/workspace/GB09_00-2010_174_Oxford_parks/delivery/20100825

comment:17 Changed 9 years ago by jb362

Delivery checking camera.

comment:18 Changed 9 years ago by jb362

Delivery check complete- everything good

comment:19 Changed 9 years ago by jb362

Delivery checking LiDAR

comment:20 follow-up: Changed 9 years ago by jb362

Looks like I can't do this since lag won't work on Fedora 12

comment:21 in reply to: ↑ 20 Changed 9 years ago by emca

Replying to jb362:

Looks like I can't do this since lag won't work on Fedora 12

You can check the files using either grass or fugroviewer as an alternative to LAG.

comment:22 Changed 9 years ago by ella

LiDAR checking complete, all looks good.

comment:23 Changed 9 years ago by emca

LiDAR and Photography dispatched to PI 26/08/2010

comment:24 Changed 9 years ago by emca

  • Description modified (diff)

comment:25 Changed 9 years ago by emca

LiDAR and Photography have been rsync'd to
/users/rsg/arsf/arsf_data/2010/flight_data/uk/GB09_00-2010_174_Oxford_Parks

comment:26 Changed 8 years ago by anch

Hawk data reprocessed and added to delivery 20101206

comment:27 Changed 8 years ago by mark1

Hyperspectral delivery checked and looks good.

comment:28 Changed 8 years ago by anch

  • Description modified (diff)

Hawk and Eagle delivered 20/12/2010

comment:29 Changed 7 years ago by knpa

  • Milestone set to 2010 data processing completion
  • Owner changed from anch to knpa
  • Status changed from accepted to assigned

Hyperspectral reprocessing was done by jaho.

Looks finished so doing a delivery check.

comment:30 Changed 7 years ago by knpa

  • Component changed from Processing: general to Archiving

Delivery check complete.

Made the following changes:

  • There were some extra incorrect mapped files present, I removed these
  • Corrected the DEM name in the example aplcorr command in the ReadMe
  • Removed the OSTN02 file as we're not supposed to distribute this

This flight does not need to be re-delivered, so this is finished.

comment:31 Changed 5 years ago by knpa

Merged and deleted workspace version.

comment:32 Changed 3 years ago by dac

Archiving

Started uploading, LiDAR, camera and raw data to NEODC.

comment:33 Changed 3 years ago by dac

  • Resolution set to fixed
  • Status changed from assigned to closed

Archiving

Archived data now available from NEODC: http://browse.ceda.ac.uk/browse/neodc/arsf/2010/GB09_00/GB09_00-2010_174_Oxford_Parks

Marking as closed.

Note: See TracTickets for help on using tickets.